# ProLLM: Protein Chain-of-Thoughts Enhanced LLM for Protein-Protein Interaction Prediction

### v1

@@ 3.2.1 Embedding Replacement
We use a new embedding mechanism facilitated by the ProTrans [7] model to perform embedding replacement. ProTrans is a large-scale protein model capable of transforming protein sequences into embedding vectors with the biophysical and structural features of proteins.
@@
We add all protein IDs in the protein dataset and their corresponding embeddings to the T5 vocabulary. Because ProtTrans is trained based on T5-large, the process of adding to the vocabulary does not involve any change in dimensions, and there is no information loss.
@@ 3.2.2 Instruction Finetuning
Mol-Instructions (Mol) [42] dataset is applied to conduct instruction fine-tuning on our model. Mol is a comprehensive instruction dataset designed for the field of biomolecules that aims to address the limitations of LLMs in biomolecular research. We use Mol to teach our model with knowledge of protein functions.
@@ 4.1 Experimental Settings
We undertake comprehensive evaluations on a trio of public Protein-Protein Interaction (PPI) datasets: Human [3], STRING [43], SHS27k, and SHS148k [44]. In these four datasets, we employ DFS dataset partitioning techniques. By prioritizing depth, DFS can effectively capture the step-by-step signal transmission and the hierarchical signal level of protein signaling pathways. Each dataset is split into training, validation, and testing sets, maintaining a proportion of 70%, 10%, and 20% respectively.
@@
Earlier studies on predicting PPI are not pre-trained, they do not have prior knowledge about the proteins. Hence, we choose SVM [45], DPPI [23], DNN-PPI [46], PIPR [44], GNN-PPI [4] and SemiGNN-PPI [47] as the baseline without pretraining. Furthermore, we explore how protein pre-training can influence the PPI, we include ProBERT [7], SM-1b [48], GearNet-Edge [49], and KeAP [50] as pre-trained baselines. MAPE-PPI [29] and InstrucGLM [9] have two versions: one with pre-training and one without pre-training.
@@
We selected the micro-F1 score [51] as our evaluation metric because the PPI dataset exhibits class imbalance, making the F1 score a very relevant reference. Note that micro-F1 is widely used in the protein-protein interaction task [52]. Each dataset follows a 70% training, 10% validation, and 20% testing data split. Subsequently, we will choose different random seeds for training and testing, conducting a total of 10 tests. The mean micro-F1 score across these trials will serve as the definitive measure of model performance, with the accompanying standard deviation reflecting variability in different experimental runs.
@@ 4.2 Comparative Experiment (RQ1)
We compare the performance of ProLLM with other baselines (w/ and w/o additional pre-training data) on four datasets in Table 1. Note that here we use Flan-T5-large [33] as the backbone model. Based on the results, we can make three important observations: (1) Our method outperforms other baselines without pre-training. Although InstructGLM is also LLM based, it lags behind ProLLM; (2) As for the models pre-trained on protein dataset, they cannot achieve the performance of ProLLM without prior knowledge; (3) ProLLM outperforms GearNet-Edge, KeAP and MAPE-PPI, although they utilize a significantly larger dataset comprised of structural and knowledge graph data for pre-training than the Mol dataset.
@@
ProLLM(Flan-T5-large) - 87.32$\pm$1.93 83.87$\pm$3.54 85.13$\pm$1.86 87.12$\pm$1.68
@@
InstructGLM Mol dataset 85.71$\pm$2.01 81.63$\pm$3.44 83.41$\pm$1.78 85.25$\pm$1.72
ProLLM(Flan-T5-large) Mol dataset 91.05$\pm$1.63 85.32$\pm$3.29 87.66$\pm$1.68 89.21$\pm$1.45
@@ 4.3 Influence of Different Backbones (RQ2)
We choose Flan-T5-base, Flan-T5-large, Flan-T5-XL [33] and LLaMA-7b [32] models as the backbone for ProLLM. We report the micro-F1 performance comparison as in Table 2. We should replace the embedding in Flan-T5 by ProtTrans and ProtTrans is a pretrained model based on Flan-T5-Large. Therefore, the embedding generated by ProtTrans will match Flan-T5-large better. Additionally, despite having more model parameters, LLaMA-v1-7b exhibits worse lower micro-F1 scores in four datasets compared to lighter models: Flan-T5-base, Flan-T5-large, and Flan-T5-XL. Furthermore, the greater standard deviation of LLaMA-v1-7b highlights its instability.
@@
ProLLM-Flan-T5-base - 82.62$\pm$2.01 76.37$\pm$4.11 78.19$\pm$1.82 80.63$\pm$1.97
ProLLM-Flan-T5-large - 87.32$\pm$1.93 83.87$\pm$3.54 85.13$\pm$1.86 87.12$\pm$1.68
ProLLM-Flan-T5-XL - 84.32$\pm$2.65 79.51$\pm$3.69 81.73$\pm$2.21 82.94$\pm$1.93
ProLLM-LLaMA-v1-7b - 81.75$\pm$4.46 74.29$\pm$6.26 77.08$\pm$3.59 79.41$\pm$2.87
ProLLM-Flan-T5-base Mol dataset 87.92$\pm$2.07 82.84$\pm$3.97 85.07$\pm$1.87 87.11$\pm$1.88
ProLLM-Flan-T5-large Mol dataset 91.05$\pm$1.63 85.32$\pm$3.29 87.66$\pm$1.93 89.21$\pm$1.45
ProLLM-Flan-T5-XL Mol dataset 89.16$\pm$2.41 83.47$\pm$3.12 86.13$\pm$1.97 87.97$\pm$1.78
ProLLM-LLaMA-v1-7b Mol dataset 87.08$\pm$4.72 82.39$\pm$6.31 84.61$\pm$3.53 86.71$\pm$2.61
@@ 4.4 Ablation Study (RQ3)
In our ablation study (Table 3), we evaluated the impact of different configurations on PPI prediction. The configurations are as follows: ProLLM w/o ProCoT: This setup shuffles ProCoT data, disrupting signaling pathways to mimic the model’s performance without pathway understanding. It limits the model’s ability to learn from signaling sequences. ProLLM relies on memorizing fixed relations instead of reasoning through intermediate protein relationships; ProLLM w/o Embedding Replacement: It evaluates the effect of protein-specific embedding features; ProLLM w/o Instruction Fine-tuning: This setup examines the model’s capability to predict protein-protein interactions without the application of instruction fine-tuning on Mol dataset; ProLLM w/o Embedding and Instruction Fine-tuning: This configuration tests the model’s performance without utilizing both protein-specific embedding features and instruction fine-tuning on Mol dataset.
In our ablation study, we have identified that the presence of Procot has the most significant impact on the performance of ProLLM. Our experiments revealed that introducing Procot led to substantial improvements in the performance of the model. Additionally, we explored other techniques such as embedding replacement and instruction fine-tuning on the Mol dataset. While these approaches did show some positive effects on the model’s performance, their impact was found to be comparatively smaller when compared to the influence of ProCoT.

### v2

@@ 3.2.1 Embedding replacement
We use a new embedding mechanism facilitated by the ProTrans (Elnaggar et al., 2021a) model to perform embedding replacement. ProTrans is a large-scale protein model capable of transforming protein sequences into embedding vectors with the biophysical and structural features of proteins.
@@
We add all protein IDs in the protein dataset and their corresponding embeddings to the T5 vocabulary to solve the OOV (out of vocabulary) problem: The T5 vocabulary does not contain words for protein ID numbers. Because ProtTrans is trained based on T5-large, the process of adding to the vocabulary does not involve any change in dimensions, and there is no information loss. Additionally, this approach can also avoid the issue where a protein ID is tokenized into multiple tokens, causing the model to fail to understand it.
@@ 3.2.2 Instruction Finetuning
Mol-Instructions (Mol) (Fang et al., 2023) dataset is applied to conduct instruction fine-tuning on our model. Mol is a comprehensive instruction dataset designed for the field of biomolecules that aims to address the limitations of LLMs in biomolecular research. We use Mol to teach our model with knowledge of protein functions.
@@ 4.1 Experimental Settings
We undertake comprehensive evaluations on a trio of public Protein-Protein Interaction (PPI) datasets: Human (Song et al., 2022), STRING (Szklarczyk et al., 2019), SHS27k, and SHS148k (Chen et al., 2019). In these four datasets, we employ DFS dataset partitioning techniques. By prioritizing depth, DFS can effectively capture the step-by-step signal transmission and the hierarchical signal level of protein signaling pathways. Each dataset is split into training, validation, and testing sets, maintaining a proportion of 70%, 10%, and 20% respectively.
@@
Earlier studies on predicting PPI are not pre-trained, they do not have prior knowledge about the proteins. Hence, we choose SVM (Romero-Molina et al., 2019), DPPI (Hashemifar et al., 2018), DNN-PPI (Li et al., 2018), PIPR (Chen et al., 2019), GNN-PPI (Lv et al., 2021) and SemiGNN-PPI (Zhao et al., 2023) as the baseline without pre-training. Furthermore, we explore how protein pre-training can influence the PPI, we include ProBERT (Elnaggar et al., 2021a), SM-1b (Rives et al., 2021), GearNet-Edge (Zhang et al., 2022), and KeAP (Zhou et al., 2022) as pre-trained baselines. MAPE-PPI (Wu et al., 2024) and InstrucGLM (Ye et al., 2024) have two versions: one with pre-training and one without pre-training.
@@
We selected the micro-F1 score (Harbecke et al., 2022) as our evaluation metric because the PPI dataset exhibits class imbalance, making the F1 score a very relevant reference. Note that micro-F1 is widely used in the protein-protein interaction task (Tran & Kavuluru, 2018). Each dataset follows a 70% training, 10% validation, and 20% testing data split. Subsequently, we will choose different random seeds for training and testing, conducting a total of 10 tests. The mean micro-F1 score across these trials will serve as the definitive measure of model performance, with the accompanying standard deviation reflecting variability in different experimental runs.
@@ 4.2 Comparative Experiment (RQ1)
We compare the performance of ProLLM with other baselines (w/ and w/o additional pre-training data) on four datasets in Table 1. Note that here we use Flan-T5-large (Raffel et al., 2020) as the backbone model. Based on the results, we can make three important observations: (1) Our method outperforms other baselines without pre-training. Although InstructGLM is also LLM based, it lags behind ProLLM; (2) As for the models pre-trained on protein dataset, they cannot achieve the performance of ProLLM without prior knowledge; (3) ProLLM outperforms GearNet-Edge, KeAP, and MAPE-PPI, although they utilize a significantly larger dataset comprised of structural and knowledge graph data for pre-training than the Mol dataset.
@@
ProLLM(Flan-T5-large) - 87.32$\pm$1.93 75.13$\pm$3.76 85.13$\pm$1.86 87.12$\pm$1.68
@@
InstructGLM Mol dataset 85.71$\pm$2.01 75.64$\pm$3.49 83.41$\pm$1.78 85.25$\pm$1.72
ProLLM(Flan-T5-large) Mol dataset 91.05$\pm$1.63 78.09$\pm$3.24 87.66$\pm$1.68 89.21$\pm$1.45
@@ 4.3 Influence of Different Backbones (RQ2)
We choose Flan-T5-base, Flan-T5-large, Flan-T5-XL (Raffel et al., 2020) and LLaMA-7b (Touvron et al., 2023) models as the backbone for ProLLM. We report the micro-F1 performance comparison as in Table 2. We should replace the embedding in Flan-T5 with ProtTrans and ProtTrans is a pre-trained model based on Flan-T5-Large. Therefore, the embedding generated by ProtTrans will match Flan-T5-large better. Additionally, despite having more model parameters, LLaMA-v1-7b exhibits worse lower micro-F1 scores in four datasets compared to lighter models: Flan-T5-base, Flan-T5-large, and Flan-T5-XL. Furthermore, the greater standard deviation of LLaMA-v1-7b highlights its instability.
@@
ProLLM-Flan-T5-base - 82.62$\pm$2.01 71.67$\pm$4.36 78.19$\pm$1.82 80.63$\pm$1.97
ProLLM-Flan-T5-large - 87.32$\pm$1.93 75.13$\pm$3.76 85.13$\pm$1.86 87.12$\pm$1.68
ProLLM-Flan-T5-XL - 84.32$\pm$2.65 73.28$\pm$3.96 81.73$\pm$2.21 82.94$\pm$1.93
ProLLM-LLaMA-v1-7b - 81.75$\pm$4.46 70.33$\pm$6.52 77.08$\pm$3.59 79.41$\pm$2.87
ProLLM-Flan-T5-base Mol dataset 87.92$\pm$2.07 74.28$\pm$4.09 85.07$\pm$1.87 87.11$\pm$1.88
ProLLM-Flan-T5-large Mol dataset 91.05$\pm$1.63 78.09$\pm$3.24 87.66$\pm$1.93 89.21$\pm$1.45
ProLLM-Flan-T5-XL Mol dataset 89.16$\pm$2.41 75.96$\pm$3.38 86.13$\pm$1.97 87.97$\pm$1.78
ProLLM-LLaMA-v1-7b Mol dataset 87.08$\pm$4.72 74.87$\pm$6.51 84.61$\pm$3.53 86.71$\pm$2.61
@@ 4.4 Ablation Study (RQ3)
In our ablation study (Table 3), we evaluated the impact of different configurations on PPI prediction. The configurations are as follows: ProLLM w/o ProCoT: This setup shuffles ProCoT data, disrupting signaling pathways to mimic the model’s performance without pathway understanding. It limits the model’s ability to learn from signaling sequences. ProLLM relies on memorizing fixed relations instead of reasoning through intermediate protein relationships; ProLLM w/o Embedding Replacement: It will compare the model with replaced embeddings to the model without expended vocabulary, which evaluates the effect of protein-specific embedding features; ProLLM w/o Instruction Fine-tuning: This setup examines the model’s capability to predict protein-protein interactions without the application of instruction fine-tuning on Mol dataset; ProLLM w/o Embedding and Instruction Fine-tuning: This configuration tests the model’s performance without utilizing both protein-specific embedding features and instruction fine-tuning on Mol dataset.
In our ablation study, we have identified that ProCoT has the most significant impact on the performance of ProLLM. Our experiments revealed that introducing ProCoT led to substantial improvements in the performance of the model. Additionally, we explored other techniques such as embedding replacement and instruction fine-tuning on the Mol dataset. While these approaches did show some positive effects on the model’s performance, their impact was found to be comparatively smaller when compared to the influence of ProCoT.
@@ 5 Conclusions and Future Work
5.1 Acknowledgement
We thank Wenyue Hua, Kai Mei and Taowen Wang for their valuable discussions and suggestions during the project.